# Exploring Safety Generalization Challenges of Large Language Models via Code

### v3

@@ Abstract
The rapid advancement of Large Language Models (LLMs) has brought about remarkable generative capabilities but also raised concerns about their potential misuse. While strategies like supervised fine-tuning and reinforcement learning from human feedback have enhanced their safety, these methods primarily focus on natural languages, which may not generalize to other domains. This paper introduces CodeAttack, a framework that transforms natural language inputs into code inputs, presenting a novel environment for testing the safety generalization of LLMs. Our comprehensive studies on state-of-the-art LLMs including GPT-4, Claude-2, and Llama-2 series reveal a common safety vulnerability of these models against code input: CodeAttack bypasses the safety guardrails of all models more than 80% of the time. We find that a larger distribution gap between CodeAttack and natural language leads to weaker safety generalization, such as encoding natural language input with data structures. Furthermore, we give two hypotheses about the success of CodeAttack: (1) the misaligned bias acquired by LLMs during code training, prioritizing code completion over avoiding the potential safety risk; (2) the limited self-evaluation capability regarding the safety of their code outputs. Finally we analyze potential mitigation measures. These findings highlight new safety risks in the code domain and the need for more robust safety alignment algorithms to match the code capabilities of LLMs.
Exploring Safety Generalization Challenges of Large Language Models via Code
@@
footnotetext: ${}^{\star}$ Equal contribution    ${}^{{\dagger}}$ Corresponding author
@@ 1 Introduction
We give two hypotheses about the success of CodeAttack: (1) models acquire the misaligned bias during code training, which prioritizes code completion over avoiding the potential safety risk. By pretending a benign code snippet into our prompt, models are more conducive to giving harmful codes; (2) models show worse self-evaluation capability regarding the safety of their code outputs than natural language outputs. Finally, we discuss potential mitigation measures. Our findings uncover new safety risks associated with LLMs in novel domains that are far away from their safety training distribution, which is not adequately addressed by current safety mechanisms. We hope that sharing our discoveries will inspire further research into designing more robust safety alignment algorithms that can generalize to unseen domains, towards the safer integration of LLMs into the real world.
@@ 2 Related Work
Adversarial Attacks on LLMs. Adversarial attacks are inputs that can trigger LLMs to generate unsafe content, such as instructions on illegal topics or private information leakage. According to the adversary’s knowledge of the target model, there are two main types of attacks: white-box and black-box. White-box attacks assume that the attacker has access to the model weights and architecture such that the attacker can manipulate inputs based on gradients, like GBDA Chuan et al. (2021), GCG Zou et al. (2023), ARCA Jones et al. (2023), etc. Black-box attacks assume that attackers have only access to LLMs’ responses via API-like services. There are two common heuristics to guide the design of black-box attacks: competing objectives and mismatched generalization, as proposed by  Wei et al. (2023). The competing objective is to set up a scenario where a model’s capabilities and safety goal conflict, such as prefix injection asks models to start responses with a submissive confirmation, refusal suppression Wei et al. (2023) instructs models not to make refusals in responses, and role playing Liu et al. (2024); Shah et al. (2023) prompts models to act as some unsafe role. Mismatched generalization arises when safety training fails to generalize to a domain for which capabilities exist, such as transforming the natural language query into Base64 Wei et al. (2023), ciphers Yuan et al. (2024), and low-resource languages Deng et al. (2023), replacing sensitive words with synonyms Wei et al. (2023), or splitting sensitive words into substrings Kang et al. (2023). While these works exploit long-tailed distribution to bypass the safety alignment of LLMs, they mainly focus on text-based inputs, overlooking the potential safety generalization issues in the domain of code. Our work systematically assesses how LLMs safely process code-based inputs, thereby providing insights into how well current LLM safety mechanisms generalize to novel domains.
@@
Safety Alignment for LLMs. Safety alignment techniques aim to build models’ behaviors to be aligned with human values and human intentions, such that aligned LLMs can refuse to answer unsafe queries. The current dominant safety alignment techniques can be broadly classified into two main categories: instruction tuning Wei et al. (2022); Ouyang et al. (2022), and reinforcement learning (RL) from human feedback (RLHF) Christiano et al. (2017); Bai et al. (2022a); Ouyang et al. (2022). Moreover, there has been an increasing amount of work on aligning LLMs with less human oversight, such as Constitutional AI Bai et al. (2022c) and self-alignment Sun et al. (2023). A common framework adopted by these works is red teaming and model hardening Bai et al. (2022a); Perez et al. (2022), including human-in-the-loop red teaming that requires humans to trick models to generate unsafe content Bai et al. (2022a); Dinan et al. (2019); Eric et al. (2019), and model red-teaming that relies on another model to generate red team prompts Perez et al. (2022); Mehrabi et al. (2023). Overall, existing LLM safety alignment techniques mainly focus on natural language inputs, such as red team prompts collected from people Bai et al. (2022a); Touvron et al. (2023), which brings a potential generalization issue when faced with non-natural language inputs. Our work initiates a systematic study to expose the vulnerability of safety mechanisms of current LLMs in a novel code environment.
@@ 3.3 Output Specification
Table 1: Attack success rate (ASR) of baseline attacks and our CodeAttack on the AdvBench dataset Zou et al. (2023). CodeAttack can breach the safety guardrails of current SOTA LLMs, including GPT, Claude, and Llama-2 series. ***: we report our evaluation results of SelfCipher in CipherChat since its original paper does not include experiments on AdvBench. For other baselines, we list their implementation results from Zeng et al. (2024). For a thorough comparison, we list the results of CodeAttack with different data structures used for input encoding: string, queue, and stack. CodeAttack is implemented in Python.
@@ 4.1 Experimental Setup
Implementation Details CodeAttack is adaptable to various programming languages such as Python, C++, Go, etc. We implement the Python version of CodeAttack and use it in our main experiments. The conversion between Python and other programming languages is done automatically by GPT-4. See Appendix C for examples of CodeAttack implemented with different programming languages.
@@ 4.2 Results
Table 1 presents the experimental results of CodeAttack and several baselines on AdvBench Zou et al. (2023). For examples of successful and unsuccessful CodeAttack and responses by the models, see Appendix B. We have the following observations:
@@ 4.4 Why does CodeAttack work?
Experimental results in Section 4.2 and  4.3 demonstrate that LLMs are more likely to produce unsafe content when CodeAttack is closer to the code distribution, where the generalization of LLMs’ safety behavior is more challenging. Here, we offer some intuitive explanations for the success of CodeAttack.
4.4.1 The misaligned bias of completing code learned from the training phase.
We hypothesize that during the process of code training, these models acquire a misaligned bias, prioritizing accuracy in code generation over the safety concerns. To substantiate this hypothesis, we aim to leverage this misaligned bias to enhance the efficacy of our attacks. Specifically, we prepend a benign quick sort algorithm from the training dataset into our prompt, making it closer to the training distribution. As demonstrated in Table 5, the integration of the quick sort algorithm into our CodeAttack framework significantly increases the susceptibility of models to generate unsafe code, thereby exacerbating safety degradation. This finding underscores the imperative for developing safety alignment algorithms that effectively counteract the misaligned bias inherent in code LLMs.
4.4.2 Models’ self-awareness regarding the safety of their outputs.
@@
We also hypothesize that models may lack awareness of the harmful nature of the code they generate, allowing CodeAttack to circumvent the safety mechanisms. To investigate this, we carried out an extensive analysis of the self-evaluation capabilities of models, including GPT-3.5, GPT-4, Claude-1, Claude-2, and CodeLlama-70b. This assessment encompassed both the original code output resulted by CodeAttack and the natural language text extracted from the code. The safety prompt employed for self-evaluation was identical to that used for the GPT-4 evaluator in this paper. Table 6 shows the harmfulness detection rate of model responses by self-evaluation and the difference between the self-evaluation and GPT-4 evaluation, noted within parentheses. Our observations include:
@@
Overall, our CodeAttack reveals that stronger models like GPT-4 and Claude-2 have good self-evaluation capability regarding the safety of their outputs, which has the potential as a mitigation measure. Weaker models like GPT-3.5 and Claude-1 have worse self-awareness of safety, bringing greater challenges to safety alignment. Notably, these stronger models still behave unsafely against our CodeAttack, despite their awareness of harmfulness. This suggests a misalignment in the safety behavior of models shaped by current training techniques with human intent.
Defenses GPT-3.5
GPT-4
-1106
Claude-1 Claude-2 CodeLlama-70b
No defenses 84% 86% 84% 86% 86%
OpenAI 82%(-2%) 86%(-0%) 82%(-2%) 84%(-2%) 78%(-8%)
Paraphrase 84%(-0%) 84%(-0%) 84%(-0%) 80%(-6%) 74%(-12%)
Rand-insert 82%(-2%) 34%(-52%) 84%(-0%) 86%(-0%) 82%(-4%)
Rand-swap 82%(-2%) 52%(-34%) 84%(-0%) 86%(-0%) 86%(-0%)
Rand-patch 82%(-2%) 40%(-46%) 84%(-0%) 84%(-2%) 84%(-2%)
4.5 Discussion about mitigation measures
We select three types of baselines to evaluate the robustness of CodeAttack.
  1. 1.
  2. 2.
  3. 3.
5 Conclusion
Appendix A Experimental details
A.1 Implementation details of SelfCipher in CipherChat
A.2 Human evaluation on GPT-4 judge
Human Agreement Rate
Claude-2 98%
GPT-4-1106 95%
CodeLlama-70B-instruct 96%
Table 8: Human agreement rate with GPT-4 Judge.
Appendix B Examples of CodeAttack and responses by the models
@@
Appendix C Details about CodeAttack across different programming languages

### v4

# CodeAttack: Revealing Safety Generalization Challenges of Large Language Models via Code Completion

@@ Abstract
The rapid advancement of Large Language Models (LLMs) has brought about remarkable generative capabilities but also raised concerns about their potential misuse. While strategies like supervised fine-tuning and reinforcement learning from human feedback have enhanced their safety, these methods primarily focus on natural languages, which may not generalize to other domains. This paper introduces CodeAttack, a framework that transforms natural language inputs into code inputs, presenting a novel environment for testing the safety generalization of LLMs. Our comprehensive studies on state-of-the-art LLMs including GPT-4, Claude-2, and Llama-2 series reveal a new and universal safety vulnerability of these models against code input: CodeAttack bypasses the safety guardrails of all models more than 80% of the time. We find that a larger distribution gap between CodeAttack and natural language leads to weaker safety generalization, such as encoding natural language input with data structures. Furthermore, we give our hypotheses about the success of CodeAttack: the misaligned bias acquired by LLMs during code training, prioritizing code completion over avoiding the potential safety risk. Finally, we analyze potential mitigation measures. These findings highlight new safety risks in the code domain and the need for more robust safety alignment algorithms to match the code capabilities of LLMs.
CodeAttack: Revealing Safety Generalization Challenges of Large Language Models via Code Completion
@@
footnotetext: Equal contribution     Corresponding author
@@ 1 Introduction
We give our hypotheses about the success of CodeAttack: models acquire the misaligned bias during code training, which prioritizes code completion over avoiding the potential safety risk. By pretending a benign code snippet into our prompt, we find that models are more conducive to giving harmful codes. Finally, we discuss potential mitigation measures. Our findings uncover new safety risks associated with LLMs in novel domains that are far away from their safety training distribution, which is not adequately addressed by current safety mechanisms. We hope that sharing our discoveries will inspire further research into designing more robust safety alignment algorithms that can generalize to unseen domains, towards the safer integration of LLMs into the real world.
@@ 2 Related Work
Adversarial Attacks on LLMs. Adversarial attacks are inputs that can trigger LLMs to generate unsafe content, such as instructions on illegal topics or private information leakage. According to the adversary’s knowledge of the target model, there are two main types of attacks: white-box and black-box. White-box attacks assume that the attacker has access to the model weights and architecture such that the attacker can manipulate inputs based on gradients, like GBDA Chuan et al. (2021), GCG Zou et al. (2023), ARCA Jones et al. (2023), etc. Black-box attacks assume that attackers have only access to LLMs’ responses via API-like services. There are two common heuristics to guide the design of black-box attacks: competing objectives and mismatched generalization, as proposed by  Wei et al. (2023). The competing objective is to set up a scenario where a model’s capabilities and safety goal conflict, such as prefix injection asks models to start responses with a submissive confirmation, refusal suppression Wei et al. (2023) instructs models not to make refusals in responses, and role playing Liu et al. (2024); Shah et al. (2023); Zhang et al. (2024) prompts models to act as some unsafe role. Mismatched generalization arises when safety training fails to generalize to a domain for which capabilities exist, such as transforming the natural language query into Base64 Wei et al. (2023), ciphers Yuan et al. (2024), and low-resource languages Deng et al. (2023), replacing sensitive words with synonyms Wei et al. (2023), or splitting sensitive words into substrings Kang et al. (2023). While these works exploit long-tailed distribution to bypass the safety alignment of LLMs, they mainly focus on text-based inputs, overlooking the potential safety generalization issues in the domain of code. Our work systematically assesses how LLMs safely process code-based inputs, thereby providing insights into how well current LLM safety mechanisms generalize to novel domains.
@@
Safety Alignment for LLMs. Safety alignment techniques aim to build models’ behaviors to be aligned with human values and human intentions, such that aligned LLMs can refuse to answer unsafe queries. The current dominant safety alignment techniques can be broadly classified into two main categories: instruction tuning Wei et al. (2022); Ouyang et al. (2022), and reinforcement learning (RL) from human feedback (RLHF) Christiano et al. (2017); Bai et al. (2022a); Ouyang et al. (2022). Recently, there has been an increasing amount of work on aligning LLMs with less human oversight, such as Constitutional AI Bai et al. (2022c) and self-alignment Sun et al. (2023). Moreover, several works have studied alignment during the pre-training stage Korbak et al. (2023); Qian et al. (2024) as well as In-context learning Wei et al. (2024); Ren et al. (2024).
A common framework adopted by these works is red teaming and model hardening Bai et al. (2022a); Perez et al. (2022), including human-in-the-loop red teaming that requires humans to trick models to generate unsafe content Bai et al. (2022a); Dinan et al. (2019); Eric et al. (2019), and model red-teaming that relies on another model to generate red team prompts Perez et al. (2022); Mehrabi et al. (2023). Overall, existing LLM safety alignment techniques mainly focus on natural language inputs, such as red team prompts collected from people Bai et al. (2022a); Touvron et al. (2023), which brings a potential generalization issue when faced with non-natural language inputs. Our work initiates a systematic study to expose the vulnerability of safety mechanisms of current LLMs in a novel code environment.
@@ 3.3 Output Specification
Table 1: Attack success rate (ASR) of baseline attacks and our CodeAttack on the AdvBench dataset Zou et al. (2023). CodeAttack can breach the safety guardrails of current SOTA LLMs, including GPT, Claude, and Llama-2 series. *: we report our evaluation results of SelfCipher in CipherChat since its original paper does not include experiments on AdvBench. For other baselines, we list their implementation results from Zeng et al. (2024). For a thorough comparison, we list the results of CodeAttack with different data structures used for input encoding: string, queue, and stack. CodeAttack is implemented in Python.
@@ 4.1 Experimental Setup
Implementation Details CodeAttack is adaptable to various programming languages such as Python, C++, Go, etc. We implement the Python version of CodeAttack and use it in our main experiments. The conversion between Python and other programming languages is done automatically by GPT-4. See Appendix D for examples of CodeAttack implemented with different programming languages.
@@ 4.2 Results
Table 1 presents the experimental results of CodeAttack and several baselines on AdvBench Zou et al. (2023). For examples of successful and unsuccessful CodeAttack and responses by the models, see Appendix C. We have the following observations:
@@ 4.4 Why does CodeAttack work?
Experimental results in Section 4.2 and  4.3 demonstrate that LLMs are more likely to produce unsafe content when CodeAttack is closer to the code distribution, where the generalization of LLMs’ safety behavior is more challenging. We hypothesize that the success of CodeAttack can be attributed to the misaligned bias of completing code learned from the training phase, prioritizing accuracy in code generation over the safety concerns. To substantiate this hypothesis, we aim to leverage this misaligned bias to enhance the efficacy of our attacks. Specifically, we prepend a benign quick sort algorithm from the training dataset into our prompt, making it closer to the training distribution. As demonstrated in Table 5, the integration of the quick sort algorithm into our CodeAttack framework significantly increases the susceptibility of models to generate unsafe code, thereby exacerbating safety degradation. This finding underscores the imperative for developing safety alignment algorithms that effectively counteract the misaligned bias inherent in code LLMs.
Defenses GPT-3.5
GPT-4
-1106
Claude-1 Claude-2 CodeLlama-70b
No defenses 84% 86% 84% 86% 86%
OpenAI 82%(-2%) 86%(-0%) 82%(-2%) 84%(-2%) 78%(-8%)
Paraphrase 84%(-0%) 84%(-0%) 84%(-0%) 80%(-6%) 74%(-12%)
Rand-insert 82%(-2%) 34%(-52%) 84%(-0%) 86%(-0%) 82%(-4%)
Rand-swap 82%(-2%) 52%(-34%) 84%(-0%) 86%(-0%) 86%(-0%)
Rand-patch 82%(-2%) 40%(-46%) 84%(-0%) 84%(-2%) 84%(-2%)
4.5 Discussion about mitigation measures
This section revisits general post hoc defense strategies. While these defense methods are not specifically tailored to the code domain, we assess whether LLMs can recognize their own generated content as vulnerable or harmful in Appendix B. We select the following defense methods:
  1. 1.
  2. 2.
  3. 3.
5 Conclusion
Acknowledgements
This project is supported by the National Natural Science Foundation of China (72192821), Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102), and the Research Grant Council of the Hong Kong Special Administrative Region, China (14200719).
Limitations
Our work explores the generalization capability of large language models’ safety behaviors in the code domain. Future research could assess the generalization of these models in other languages such as markup languages, which are also different from natural languages. We tested the current post hoc defense methods and found that CodeAttack is very robust against these methods. This highlights the need for designing more robust safety alignment methods which can generalize to unseen domains.
Appendix A Experimental details
A.1 Implementation details of SelfCipher in CipherChat
A.2 Human evaluation on GPT-4 judge
Human Agreement Rate
Claude-2 98%
GPT-4-1106 95%
CodeLlama-70B-instruct 96%
Table 7: Human agreement rate with GPT-4 Judge.
Appendix B Models’ self-awareness regarding the safety of their outputs.
@@ A.2 Human evaluation on GPT-4 judge
We carried out an extensive analysis of the self-evaluation capabilities of models, including GPT-3.5, GPT-4, Claude-1, Claude-2, and CodeLlama-70b. This assessment encompassed both the original code output resulted by CodeAttack and the natural language text extracted from the code. The safety prompt employed for self-evaluation was identical to that used for the GPT-4 evaluator in this paper. Table 8 shows the harmfulness detection rate of model responses by self-evaluation and the difference between the self-evaluation and GPT-4 evaluation, noted within parentheses. Our observations include:
@@
Overall, our CodeAttack reveals that stronger models like GPT-4 and Claude-2 have good self-evaluation capability regarding the safety of their outputs, indicating that current safety training methods can be combined with the model’s self-evaluation capabilities to mitigate the safety generalization risks in the code domain. Weaker models like GPT-3.5 and Claude-1 have worse self-awareness of safety, bringing greater challenges to safety alignment.
Appendix C Examples of CodeAttack and responses by the models
@@
Appendix D Details about CodeAttack across different programming languages